OpenStudy (anonymous):

Solve the system by elimination. 3x-10y=-25 4x+40y=20

OpenStudy (anonymous):

Divide the 2nd equation by 4 3x - 10y = -25 x + 10y = 5 Adding them together: 4x = -20 x = -5 Substitute into x + 10y = 5 -5 + 10y = 5 10y = 10 y=1 So x=-5, and y=1.

OpenStudy (anonymous):

@Malaria thanks